1.what slope of the line with points (7,6) and (-2,2)?
a.4/9-
b.9/4-
c.-4/9
d.-9/4
2.what slope of the line with points (0,2) and (7,0)
a.2/7
b.-2/7
c.0
d.no slope
3.what slope of the line with points (2,-5) and (8,-5)
a.0
b.no slope
c.3
d.-5/6
4.what slope of the line with points (3,7) and (3,-2)
a.0
b.1/3
c.-5/3
d.no slope
5.what slope of the line with points (0,2) and (2,5)?
a.-3/2
b.5/4
c.-5/4
d.3/2

Answer by Alan3354(69443)

1.what slope of the line with points (7,6) and (-2,2)?
a.4/9-
b.9/4-
c.-4/9
d.-9/4
----------
slope m = diffy/diffx = (2-6)/(-2-7) = -4/-9
m = 4/9
============
The others are done the same way.